OpenStudy (anonymous):

The perimeter of a rectangle is 20 cm and its dimensions are tripled. What is the perimeter of the larger rectangle?60 ?

OpenStudy (anonymous):

Let a,b be the two sides of the rectangle. Then 2a+2b=20 cm. Now, let a' and b' be the tripled a and b, respectively: a'=3a, b'=3b. We are asked for 2a'+2b': 2a'+2b'=2.3a+2.3b=3.2a+3.2b=3(2a+2b)=60cm. You are correct. Try to figure out how the area changes.